How many real number solutions are their in this equation
X^2 + 2x - 6=0
How did you get the answer

Answer by Alan3354(69443)

How many real number solutions are their in this equation
X^2 + 2x - 6=0
-----------
Check the discriminant.
Disc = b^2 - 4ac = 2^2 - 4*1*(-6) = 28
Disc > 0 --> 2 different real solutions.